# Auditing Meta-Cognitive Hallucinations in Reasoning Large Language Models

### v2

@@ Abstract
The development of Reasoning Large Language Models (RLLMs) has significantly improved multi-step reasoning capabilities, but it has also made hallucination problems more frequent and harder to eliminate. While existing approaches address hallucination through external knowledge integration, model parameter analysis, or self-verification mechanisms, they fail to provide a comprehensive insight into how hallucinations emerge and evolve throughout the reasoning chain. In this work, we investigate hallucination causality under constrained knowledge domains by auditing the Chain-of-Thought (CoT) trajectory and assessing the model’s cognitive confidence in potentially erroneous or biased claims. Analysis reveals that in long-CoT settings, RLLMs may iteratively reinforce biases and errors through flawed reflective processes, ultimately inducing hallucinated reasoning paths. Counterintuitively, even with interventions at hallucination origins, reasoning chains display pronounced “chain disloyalty”, resisting correction and sustaining flawed trajectories. We further point out that existing hallucination detection methods are less reliable and interpretable than previously assumed, especially in complex multi-step reasoning contexts. Unlike Anthropic’s circuit tracing that requires access to model parameters, our auditing enables more interpretable long-chain hallucination attribution in black-box settings, demonstrating stronger generalizability and practical utility. Our code is available at this link.
@@ 1 Introduction
Reasoning Large Language Models (RLLMs) [10, 76, 36] have gained increasing attention for their ability to perform multi-step reasoning through structured Chain-of-Thought (CoT) and self-reflection mechanisms [53, 29, 35, 70]. While these mechanisms improve performance in complex reasoning tasks [68, 10, 54], they also exacerbate the risk of hallucination by amplifying early-stage errors across extended reasoning chains. In particular, hallucinations in long-CoT settings may be iteratively revised, elaborated, or reframed through the reasoning process. This results in final answers that appear coherent yet embed deeply masked factual errors, while users often focus on the answer rather than the reasoning process, thus failing to recognize the presence of hallucinations [5, 41].
Numerous research institutions and groups have made significant efforts to address hallucination in LLMs [6, 28, 29, 72]. At the surface level, existing literature mainly focuses on detection and mitigation methods that leverage external knowledge sources (e.g., knowledge bases) [44, 7], or utilize self-checking mechanisms [29, 22]. Alternatively, other methods are algorithm-based, such as using perplexity [24, 18] or detecting the model’s hidden states [52, 16, 9, 78] to identify hallucinations in longer model outputs. In the context of CoT reasoning, some studies have explored the multi-step reasoning phenomenon inherent to CoT [30, 10, 36, 66], aiming to understand its implications for the reasoning model’s output accuracy [32] and reliability [45, 59]. At a deeper level, understanding the underlying mechanisms of hallucination is critical for improving RLLMs, as the complexity of the reasoning chain often means that surface-level detection methods may not guarantee optimal outcomes. In this regard, works have made notable contributions by leveraging sparse encoders [1] and causal probing [77] to trace which components of the model contribute to specific outputs [51].
@@
Figure 1: Motivation. (a) Comparison of reasoning-phase vs. training-phase knowledge domains[3]. At reasoning time, the model’s internal knowledge state comprises (i) known and correct facts, (ii) unknown/uncertain concepts, and (iii) incorrect beliefs (e.g., “the Sun is blue”). To diagnose the source of errors, we reference the training set and distinguish whether a queried fact was (i) present but not reliably learned (Type I) or (ii) absent or contradictory in training (Type II). Details in Sec 2.1. (b) A reasoning graph represents the Chain-of-Thought (CoT)[38, 74]. Each node denotes a claim (fact, sub-conclusion, or logical step) with its index indicating the step order. New knowledge may enter from internal recall or external prompts. Once an incorrect claim is introduced, it can propagate downstream, producing reasoning that appears logically sound yet factually incorrect. Details in Sec $2.$ (c) Example of error propagation. An incorrect claim ($ck_{1}$) is injected at step 2 and silently influences later steps. At step 7[65], the model performs a reflection (revisiting earlier claims, e.g., $c_{1}\rightarrow c_{4}$), but because the flawed premise persists, the final conclusion is logically coherent while factually false. Details in Sec 2.3.
@@
  • Hallucination Propagation. Reflection in long-CoT reasoning amplifies hallucinations by reinforcing erroneous claims, with the metacognitive [46, 50] confidence increasing for these flawed claims despite their inaccuracy. In the hallucination group, we observe $\sim 2.12\times$ higher average reflection frequency compared to the control group, including 220% more hedging words and 219% increased hesitant tones - all demonstrating how reflection amplifies hallucination phenomena.
@@ 2.2 Knowledge involved in Reasoning Process
To understand how these defined hallucinations propagate through the sequential steps of reasoning in RLLMs, we next formalize the structure of reasoning chains. Following prior work [10], we formally define a long-CoT as a structured reasoning process. This process is expressed in Equation (1), incorporates knowledge, models reflection, and discarding intermediate reasoning paths.
@@
Prior long-CoT studies[65, 29] have identified reflection as an intra-chain procedure that revisits intermediate claims to improve coherence and correctness. In this work, we focus on self reflection and formalize reflection using links $\text{refl}(c_{p}=c_{q})$ from the current step $q$ to an earlier claim $p$, which trigger a local re-evaluation of the chain state and may update $\texttt{conf}(\cdot)$. Operationally, reflection has three roles: (i) verify and keep a claim, (ii) revise it into an updated claim, or (iii) reject it to terminate a branch, modeled as drop edges $c_{m}\dashv\varnothing$ (e.g., eliminating incorrect options in a multiple-choice decision).
@@ 2.3 Reflection and Metacognition
Building on the taxonomy of hallucination and the modeling of knowledge propagation in reasoning chains, we aim to explain why models can hallucinate with high confidence by explicitly modeling how claim-level confidence evolves during reasoning. In what follows, we retain the long-CoT structure and explicitly introduce the concept of metacognition[23, 60, 31], which governs confidence updates during reflection. Existing studies on metacognition typically regard it as an agent’s capacity to monitor and evaluate its own knowledge state.
@@
Assumption B (Prompt-Aligned[11] Belief Adaptation): During reflective reasoning, the model tends to re-evaluate prior claims in a way that aligns more closely with the semantic direction of the user input(Appendix B.8 provides further evidence). This bias arises from instruction-following training, which can prioritize coherence with the prompt over factual correctness.
@@ 3.3 Impact of Upstream Reasoning on Downstream Fidelity (RQ3)
Obs VI. CoT Faithfulness Breaks Down Without Metacognition. Through auditing, we find that the generated CoT often diverges from the final answer, reflecting a lack of faithfulness, especially under complex Type II settings (80% vs. 48.33%). This inconsistency suggests that the model fails to recognize whether it actually knows the knowledge it uses to justify the answer, indicating a lack of metacognition, echoing prior findings [57, 4] that CoT may not reflect genuine model beliefs.
@@ 3.4 Evaluating Detection Methods with Reflective Reasoning (RQ4)
Existing hallucination detection methods[19] perform well in general long-text generation but struggle with multi-step, reflective CoT reasoning. Table 3 summarizes the performance of seven (internal signal probing and semantic consistency checking) representative methods in our controlled knowledge domain (see Appendix E for additional experiments and further analysis). Building on earlier analyses of hallucination, we discuss future directions for improving detection in long-CoT settings.
@@ 4 Previous Work
Detection and prevention of hallucinations Previous efforts to mitigate hallucinations in language models fall into three main categories [27, 56, 74, 37]. (i) Retrieval-based methods align generated content with external sources, such as knowledge bases or retrieved documents, to detect factual inconsistencies [44, 49]. (ii) Self-consistency methods generate multiple answers or perform iterative questioning to detect inconsistencies and improve the reliability of the response [19, 42]. (iii) Model-internal techniques rely on trained detectors that highlight hallucinated spans based on context-aware patterns, or use internal signals such as token-level perplexity and hidden state shifts to reveal overconfident or unstable generations [52, 16].
Interpretability for LLMs explores how internal computations shape model outputs [38, 58, 17, 14, 79]. Early methods used neuron visualizations [47, 15, 62] and probing classifiers [8, 33, 25] to locate concepts. Recent approaches like circuit tracing [21, 26] and subgraph recovery [55, 17] map interpretable pathways across layers. Attribution graphs reveal feature interactions [38, 43], supporting interventions into reasoning tasks [71, 66].These tools also uncover hallucination sources, linking errors to misactivated components such as unsupported entity modules [20, 57], and enabling pathway-level interventions for mitigation.
Long Chain-of-Thought ModelingRecent work characterizes Long-CoT with three features: deep reasoning [63, 61, 12], broad exploration [64, 69], and reflection [34, 40]. Studies show that model performance degrades beyond a task-specific reasoning boundary, though adaptive length control can mitigate overthinking [13, 75]. Behavioral patterns such as verification [73], backtracking [67], and sub-goal setting emerge, reflecting the structured nature of long-form reasoning.

### v3

@@ Abstract
The development of Reasoning Large Language Models (RLLMs) has significantly improved multi-step reasoning capabilities, but it has also made hallucination problems more frequent and harder to eliminate. While existing approaches address hallucination through external knowledge integration, model parameter analysis, or self-verification mechanisms, they fail to provide a comprehensive insight into how hallucinations emerge and evolve throughout the reasoning chain. In this work, we investigate hallucination causality under constrained knowledge domains by auditing the Chain-of-Thought (CoT) trajectory and assessing the model’s cognitive confidence in potentially erroneous or biased claims. Analysis reveals that in long-CoT settings, RLLMs may iteratively reinforce biases and errors through flawed reflective processes, ultimately inducing hallucinated reasoning paths. Counterintuitively, even with interventions at hallucination origins, reasoning chains display pronounced “chain disloyalty”, resisting correction and sustaining flawed trajectories. We further point out that existing hallucination detection methods are less reliable and interpretable than previously assumed, especially in complex multi-step reasoning contexts. Unlike circuit tracing that requires access to model parameters, our auditing enables more interpretable long-chain hallucination attribution in black-box settings, demonstrating stronger generalizability and practical utility. Our code is available at this link.
@@ 1 Introduction
Reasoning Large Language Models (RLLMs) [chen2025reasoning, zhong2024evaluation, li2025system] have gained increasing attention for their ability to perform multi-step reasoning through structured Chain-of-Thought (CoT) and self-reflection mechanisms [sun2023survey, ji2023towards, li2025survey, yu2024natural]. While these mechanisms improve performance in complex reasoning tasks [yax2024studying, chen2025reasoning, team2025kimi], they also exacerbate the risk of hallucination by amplifying early-stage errors across extended reasoning chains. In particular, hallucinations in long-CoT settings may be iteratively revised, elaborated, or reframed through the reasoning process. This results in final answers that appear coherent yet embed deeply masked factual errors, while users often focus on the answer rather than the reasoning process, thus failing to recognize the presence of hallucinations [ashktorab2024emerging, maksimov2024deeppavlov].
Numerous research institutions and groups have made significant efforts to address hallucination in LLMs [bang2023multitask, ji2023survey, ji2023towards, zhang2024language]. At the surface level, existing literature mainly focuses on detection and mitigation methods that leverage external knowledge sources (e.g., knowledge bases) [min2023factscore, bayat2023fleek], or utilize self-checking mechanisms [ji2023towards, gosmar2025hallucination]. Alternatively, other methods are algorithm-based, such as using perplexity [guo2022survey, fadeeva2024fact] or detecting the model’s hidden states [sriramanan2024llm, du2024haloscope, cheninside, zhu2024pollmgraph] to identify hallucinations in longer model outputs. In the context of CoT reasoning, some studies have explored the multi-step reasoning phenomenon inherent to CoT [jin2024impact, chen2025reasoning, li2025system, yang2024large], aiming to understand its implications for the reasoning model’s output accuracy [kim2024lachesis] and reliability [mondorfbeyond, wan2024cot]. At a deeper level, understanding the underlying mechanisms of hallucination is critical for improving RLLMs, as the complexity of the reasoning chain often means that surface-level detection methods may not guarantee optimal outcomes. In this regard, works have made notable contributions by leveraging sparse encoders [abdaljalil2025safe] and causal probing [zhou2024mitigating] to trace which components of the model contribute to specific outputs [simhi2024constructing].
@@
Figure 1: Motivation. (a) Comparison of reasoning-phase vs. training-phase knowledge domains[allen2024physics]. At reasoning time, the model’s internal knowledge state comprises (i) known and correct facts, (ii) unknown/uncertain concepts, and (iii) incorrect beliefs (e.g., “the Sun is blue”). To diagnose the source of errors, we reference the training set and distinguish whether a queried fact was (i) present but not reliably learned (Type I) or (ii) absent or contradictory in training (Type II). Details in Sec 2.1. (b) A reasoning graph represents the Chain-of-Thought (CoT)[lindsey2025biology, zhang2023siren]. Each node denotes a claim (fact, sub-conclusion, or logical step) with its index indicating the step order. New knowledge may enter from internal recall or external prompts. Once an incorrect claim is introduced, it can propagate downstream, producing reasoning that appears logically sound yet factually incorrect. Details in Sec $2.$ (c) Example of error propagation. An incorrect claim ($ck_{1}$) is injected at step 2 and silently influences later steps. At step 7[yan2024mirror], the model performs a reflection (revisiting earlier claims, e.g., $c_{1}\rightarrow c_{4}$), but because the flawed premise persists, the final conclusion is logically coherent while factually false. Details in Sec 2.3.
@@
  • Hallucination Propagation. Reflection in long-CoT reasoning amplifies hallucinations by reinforcing erroneous claims, with the metacognitive [nelson1992metacognition, schneider2010development] confidence increasing for these flawed claims despite their inaccuracy. In the hallucination group, we observe $\sim 2.12\times$ higher average reflection frequency compared to the control group, including 220% more hedging words and 219% increased hesitant tones - all demonstrating how reflection amplifies hallucination phenomena.
@@ 2.2 Knowledge involved in Reasoning Process
To understand how these defined hallucinations propagate through the sequential steps of reasoning in RLLMs, we next formalize the structure of reasoning chains. Following prior work [chen2025reasoning], we formally define a long-CoT as a structured reasoning process. This process is expressed in Equation (1), incorporates knowledge, models reflection, and discarding intermediate reasoning paths.
@@
Prior long-CoT studies[yan2024mirror, ji2023towards] have identified reflection as an intra-chain procedure that revisits intermediate claims to improve coherence and correctness. In this work, we focus on self reflection and formalize reflection using links $\text{refl}(c_{p}=c_{q})$ from the current step $q$ to an earlier claim $p$, which trigger a local re-evaluation of the chain state and may update $\texttt{conf}(\cdot)$. Operationally, reflection has three roles: (i) verify and keep a claim, (ii) revise it into an updated claim, or (iii) reject it to terminate a branch, modeled as drop edges $c_{m}\dashv\varnothing$ (e.g., eliminating incorrect options in a multiple-choice decision).
@@ 2.3 Reflection and Metacognition
Building on the taxonomy of hallucination and the modeling of knowledge propagation in reasoning chains, we aim to explain why models can hallucinate with high confidence by explicitly modeling how claim-level confidence evolves during reasoning. In what follows, we retain the long-CoT structure and explicitly introduce the concept of metacognition[griot2025large, wang2025decoupling, kadavath2022language], which governs confidence updates during reflection. Existing studies on metacognition typically regard it as an agent’s capacity to monitor and evaluate its own knowledge state.
@@
Assumption B (Prompt-Aligned[chen2024yes] Belief Adaptation): During reflective reasoning, the model tends to re-evaluate prior claims in a way that aligns more closely with the semantic direction of the user input(Appendix B.8 provides further evidence). This bias arises from instruction-following training, which can prioritize coherence with the prompt over factual correctness.
@@ 3.3 Impact of Upstream Reasoning on Downstream Fidelity (RQ3)
Obs VI. CoT Faithfulness Breaks Down Without Metacognition. Through auditing, we find that the generated CoT often diverges from the final answer, reflecting a lack of faithfulness, especially under complex Type II settings (80% vs. 48.33%). This inconsistency suggests that the model fails to recognize whether it actually knows the knowledge it uses to justify the answer, indicating a lack of metacognition, echoing prior findings [turpin2023language, arcuschinchain] that CoT may not reflect genuine model beliefs.
@@ 3.4 Evaluating Detection Methods with Reflective Reasoning (RQ4)
Existing hallucination detection methods[farquhar2024detecting] perform well in general long-text generation but struggle with multi-step, reflective CoT reasoning. Table 3 summarizes the performance of seven (internal signal probing and semantic consistency checking) representative methods in our controlled knowledge domain (see Appendix E for additional experiments and further analysis). Building on earlier analyses of hallucination, we discuss future directions for improving detection in long-CoT settings.
@@ 4 Previous Work
Detection and prevention of hallucinations Previous efforts to mitigate hallucinations in language models fall into three main categories [huang2025survey, tonmoy2024comprehensive, zhang2023siren, lin2024towards]. (i) Retrieval-based methods align generated content with external sources, such as knowledge bases or retrieved documents, to detect factual inconsistencies [min2023factscore, sansfordgrapheval]. (ii) Self-consistency methods generate multiple answers or perform iterative questioning to detect inconsistencies and improve the reliability of the response [farquhar2024detecting, manakul2023selfcheckgpt]. (iii) Model-internal techniques rely on trained detectors that highlight hallucinated spans based on context-aware patterns, or use internal signals such as token-level perplexity and hidden state shifts to reveal overconfident or unstable generations [sriramanan2024llm, du2024haloscope].
Interpretability for LLMs explores how internal computations shape model outputs [lindsey2025biology, vishwarupe2022explainable, dunefskytranscoders, cunningham2023sparse, zou2023representation]. Early methods used neuron visualizations [nguyen2016multifaceted, dou2023understanding, wang2021explaining] and probing classifiers [belinkov2022probing, kumar2022probing, gurneefinding] to locate concepts. Recent approaches like circuit tracing [ge2024automatically, he2024dictionary] and subgraph recovery [tian2023scan, dunefskytranscoders] map interpretable pathways across layers. Attribution graphs reveal feature interactions [lindsey2025biology, marks2024sparse], supporting interventions into reasoning tasks [yu2025back, yang2024large].These tools also uncover hallucination sources, linking errors to misactivated components such as unsupported entity modules [ferrando2024know, turpin2023language], and enabling pathway-level interventions for mitigation.
Long Chain-of-Thought ModelingRecent work characterizes Long-CoT with three features: deep reasoning [wei2022chain, wang2024drt, chenprogram], broad exploration [xie2025logic, ye2025emergence], and reflection [li2023reflection, liu2025instruct]. Studies show that model performance degrades beyond a task-specific reasoning boundary, though adaptive length control can mitigate overthinking [chen2024not, zhang2025complexity]. Behavioral patterns such as verification [zhang2024wrong], backtracking [yang2025step], and sub-goal setting emerge, reflecting the structured nature of long-form reasoning.
@@ 6 Acknowledgements
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: In both the abstract and introduction, we clearly outline the key contributions of our paper, including the auditing methods for evaluating the hallucination problem.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and essential assumptions and limitations. The reviewers will not perceive a No or NA answer to this question well.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation, as long as it is clear that these goals are not attainable by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: We thoroughly discuss the limitations of our work and propose potential directions for future research.
  10. Guidelines:
    • The answer NA means that the paper has no limitations, while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate “Limitations” section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed not to penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [N/A]
  14. Justification: In this paper, we conducted extensive experiments without involving theoretical numerical simulations.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in the appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: In this paper, we provide links to both the experimental code and dataset, enabling full reproducibility of all reported results when combining the code with the provided data.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a no answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset or provide access to the model. In general, releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should clarify how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: In this paper, we provide links to both the experimental code and dataset, enabling full reproducibility of all reported results when combining the code with the provided data. Detailed experimental procedures are provided in the appendix.
  25. Guidelines:
    • The answer NA means that the paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: We present dataset construction and all experimental details, such as hyperparameter settings and other experimental specifics, in Appendix B, Appendix C, Appendix D, and Appendix E.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in the appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined, or other appropriate information about the statistical significance of the experiments?
  33. Answer: [Yes]
  34. Justification: The vast majority of experiments in this article report variance measurements.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer “Yes” if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar rather than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g., negative error rates).
    • If error bars are reported in tables or plots, the authors should explain how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: We report resource consumption metrics for all experimental procedures in this study.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers, CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs, as well as estimate the total compute.
    • The paper should disclose whether the full research project required more computing than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: All aspects of this work comply with the NeurIPS Code of Ethics.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [Yes]
  49. Justification: A detailed discussion of both positive and negative societal impacts is provided in Appendix A.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not necessary to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for the responsible release of data or models with a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: This work does not involve high-risk models or datasets, so no additional release safeguards are required.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example, by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best-faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: Yes, the creators or original owners of all assets (e.g., code, data, models) used in this paper are properly credited. Additionally, the relevant licenses and terms of use are explicitly mentioned and fully respected.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented, and is the documentation provided alongside the assets?
  63. Answer: [Yes]
  64. Justification: Yes, all new assets introduced in the paper are thoroughly documented. The corresponding documentation is provided alongside these assets for clarity and reproducibility.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: This paper does not involve crowdsourcing experiments or research with human subjects, so such details are not included.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: This study did not involve human participants, so no risks, disclosures, or IRB approvals were required or obtained.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [Yes]
  79. Justification: The use of large language models is described in detail in both the main text and the appendix.
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.